OpenStudy (anonymous):

Use the distributive property to rewrite the expression 3(x + 5) without parentheses. A. 3x + 5 B. 3x + 8 C. 3 + 5x D. 3x + 15

OpenStudy (texaschic101):

3(x + 5) = 3 * x + 3 * 5 = 3x = 15

OpenStudy (texaschic101):

oops...3x + 15

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

can u help with another one?

OpenStudy (texaschic101):

anytime :) I can help with more

OpenStudy (anonymous):

Consider this equation: 4x – 5 + 6x = 6 Which is an equivalent equation after combining like terms? A. -1 + 6x = 6 B. 10x - 5 = 6 C. 4x + 1 = 6 D. 4x - 5 = 0

OpenStudy (texaschic101):

4x - 5 + 6x = 6 10x - 5 = 6

OpenStudy (texaschic101):

you see, when adding the 4x + 6x, that gave us 10x

OpenStudy (anonymous):

makes more sense thanks you're a lot of help